OpenStudy (anonymous):

A furniture company sells 3000 chairs per week at a price of $150 apiece; when the price is lowered to $125, the company sells 3500 chairs per week. If the company wishes to liquidate its entire remaining inventory of 4250 chairs in the next week, what price should they seek?

OpenStudy (anonymous):

First, find the an equation that expresses this situation. Let's assume this relationship is a linear one. The independent variable (x) is the price of the chair and the dependent variable is the number of units sold. Find the slope, which is the change in y divided by the change in x, or \[m=\frac{y_2-y_1}{x_2-x_1}\]\[m=\frac{3500-3000}{125-150}=-20\] Put this into the slope-intercept equation: \[y=mx+b\]\[y=-20x+b\] Choose one of the points (150, 3000) or (125, 3500). Let's use the first point, (150, 3000). In this point, 150 is the x-coordinate, so we will substitute 150 for x. 3000 is the y-coordinate, so we will substitute 3000 for y. \[3000=-20(150)+b\]Solve for b: \[3000=-3000+b\]\[b=6000\]So our equation is now \[y=-20x+6000\] The question asks how much should be charged in order to sell 4250 chairs, so we substitute that for y and solve for x: \[4250=-20x+6000\]\[-1750=-20x\]\[x=87.5\]So the vendor should sell each chair for $87.50

OpenStudy (anonymous):

wow that is not what I was expecting

OpenStudy (anonymous):

Hey Can SomeOne Help Me After This One :D http://openstudy.com/groups/mathematics#/groups/mathematics/updates/4e24e52b0b8b3d38d3b761da